OpenStudy (anonymous):

is there an oblique asymptote here? f(x)=x^2-6x+4 divided by x+5

OpenStudy (anonymous):

asymptotes are not my friends....:(

OpenStudy (rock_mit182):

I would explain this to you but with limits so i don't think is a goog idea lol

OpenStudy (anonymous):

understood rock_mit182, math is quite complex

OpenStudy (anonymous):

\[f(x)=\frac{ x^2-6x+4 }{ x+5 }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

write the general form equation first which is y=mx+n

OpenStudy (anonymous):

I can do algebra with ease, the precalc is well not wanting to be friends! ;)

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

then first find m for -infinity

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

m is by definition: \[\lim_{x \rightarrow \infty}\frac{ f(x) }{ x }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so you've got\[\lim_{x \rightarrow \infty}\frac{ x^2-6x+4 }{ x^2+5x }\]

OpenStudy (anonymous):

the coefficients of x^2 are equal so this limits tends to 1 and it's the same for +infinity

OpenStudy (anonymous):

so m=1

OpenStudy (anonymous):

\[n=\lim_{x \rightarrow \infty}f(x)-mx\]

OpenStudy (anonymous):

but in the denominator it is only x+5

OpenStudy (anonymous):

pay attention i've multiplied directly with x.... [f(x)/x]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

as I've said m=1 now we need to find only n

OpenStudy (anonymous):

\[n=\lim_{x \rightarrow \infty}\frac{ x^2-6x+4 }{ x+5 }-x\]

OpenStudy (anonymous):

\[n=\lim_{x \rightarrow \infty}\frac{ x^2-6x+4-x^2-5x }{ x+5 }\]

OpenStudy (anonymous):

\[n=\lim_{x \rightarrow \infty}\frac{ -11x+4 }{ x+5 }=-11\]

OpenStudy (anonymous):

i was reading some notes the teacher gave us on oblique asymptotes and see what you were talking about the m and n

OpenStudy (anonymous):

so y=mx+n where m=1 and n=-11 so.. can you tell me the final answer and if you understood my explanation

OpenStudy (anonymous):

y=x-11

OpenStudy (anonymous):

you are a strict teacher, I like that! ;)

OpenStudy (anonymous):

I understood what you were saying...thanks a bunch and a well deserved medal is yours!

OpenStudy (mathmale):

Recognize that if the numerator is of one degree higher than that of the denominator, you're going to have a slant asymptote. Using long division, divide the denom. into the num. and you'll immediately have your slant asymptote. Label it: y = x - 11.